OpenStudy (anonymous):

(7a)^(5/8) b^(3/8). Show work. Express using a radical.

OpenStudy (sasogeek):

ummm, did you read the notes in the link I gave you earlier...? :/

OpenStudy (anonymous):

yes, and i did not quite understand them.

OpenStudy (sasogeek):

ok so in a general case, \(\huge \sqrt[n]{a^x}=a^\frac{x}{n}\) can you use this as a model to answer your questions...? \(\huge (7a)^\frac{5}{8}\cdot (b)^\frac{3}{8}= ?\)

OpenStudy (anonymous):

yes i can, but how do you know what number is x, n, and a. I don't know the substitution for the problem.

OpenStudy (sasogeek):

a = base x= exponent of base, or numerator of rational exponent. n= index, or the denominator of the rational exponent. :)

OpenStudy (sasogeek):

in this case, a= 7a and b x=5 and 3 n=8 and 8